# T-FAKE: Synthesizing Thermal Images for Facial Landmarking

### v2

@@ 3.2 The Dataset
Having trained $T_{\theta}$, we use it to thermalize the FAKE dataset which results in the thermal dataset T-FAKE. For each thermal condition, the dataset contains 100k images featuring at least 2k unique subjects. Each image is available in a simulated ‘cold’ and a ‘warm’ environment adding up to 200k images. This number of frames and subjects surpasses the numbers in other thermal landmarking datasets [34, 63, 13, 52, 37, 38, 3] known to the authors. In contrast to thermal landmarking datasets with a fixed distance between the subjects and the camera [34, 52, 38], the distance varies within an estimated range of 80cm to 150cm for the T-FAKE dataset. Moreover, the synthesized images are based on facial scans with a wide range of age and ethnicity and a balanced number of people identifying as male or female [61]. In addition to the original ground truth sparse 3DA-2D landmark annotations, we add dense 2D annotations by transferring Mediapipe [40] predictions for the FAKE to the T-FAKE dataset. This makes our dataset the first densely annotated thermal landmarking dataset. While we found many large thermal landmarking datasets to be non-public [63, 13, 52], our complete dataset with both thermal conditions and annotations is available for download.
In combination with the original FAKE dataset, we can train multimodal facial landmarkers jointly on RGB and thermal images as described in the next section. Table 1 displays a comparison between thermal landmarking datasets.
@@ 4 Facial Landmarking
Learning with synthetic images has proven to be very useful in facial landmarking [63, 61, 62]. Based on the proposed T-FAKE and the original FAKE dataset, we can train a supervised thermal or multimodal landmark prediction model for thermal and RGB images using synthetic data only. Moreover, we adapt a model to new annotation conventions by training a small-scale label adaptation model. The data flow for these models is visualized in Fig. 3.
@@ 4.1 Landmark Prediction
the authors of [62] proposed to learn the network $T_{\psi}:=(T_{\psi}^{1},T_{\psi}^{2})$ with $X\mapsto(\boldsymbol{\mu},\boldsymbol{\sigma}^{2})$, where $\boldsymbol{\mu}=(\mu_{l})_{l=1}^{L}\in(\mathbb{R}^{2})^{L}$ and $\boldsymbol{\sigma}^{2}=(\sigma_{l}^{2})_{l=1}^{L}\in\mathbb{R}^{L}$ by minimizing the loss function
$\mathcal{L}(\psi)$ $:=\sum_{l=1}^{L}\Big{(}\sum_{j=1}^{J}\log T^{2}_{\psi,l}(X^{j})+\frac{\|T^{1}_{\psi,l}(X^{j})-y_{l}^{j}\|_{2}^{2}}{2\,T^{2}_{\psi,l}(X^{j})}\Big{)}.$ (10)
During training, we clip $T^{2}_{\psi,l}\geq\varepsilon>0$ to avoid numerical instability due to too small variances. The first sum controls the location accuracy and the second one the uncertainty prediction. Having learned the network $T_{\psi}$, it produces not only the expected values $\boldsymbol{\mu}$ of the corresponding landmarks, but also a measure for its uncertainty $\boldsymbol{\sigma}^{2}$.
@@ 4.2 Label Adaptation
A common approach that is especially crucial for synthetic trained data and comparison of 2D and 3DA-2D landmarks is label adaptation [61, 62]. It allows us to adapt to different landmark conventions between the training dataset and an evaluation dataset. Given two sequences $(\hat{\mathbf{y}}_{k},\mathbf{y}_{k})_{k=1}^{K}$ with predicted landmarks $\hat{\mathbf{y}}_{k}\coloneqq(\hat{y}^{l}_{k})_{l=1}^{\hat{L}}\in\mathbb{R}^{2\hat{L}}$, and with ground truth landmarks $\mathbf{y}_{k}:=(y^{l}_{k})_{l=1}^{L}\in\mathbb{R}^{2\hat{L}}$ with possibly $\hat{L}\neq L$, we aim to train a third label adaptation model $T_{\zeta}:\mathbb{R}^{2\hat{L}}\to\mathbb{R}^{2{L}}$. Here, we use the loss
@@ 5.1 Implementation Details
Thermalization. We train on the RGB+Thermal SEJONG dataset recorded in lab conditions and the large-scale synthetic FAKE dataset featuring in-the-wild RGB images. We train on a resolution of $256\times 256$.
The resolution-agnostic U-net allows to apply our two models trained on $256\times 256$ images to the FAKE dataset at the original $512\times 512$ resolution to generate the T-FAKE dataset. This results in the up-to-date largest thermal landmarking dataset with 200k images of at least 2k unique subjects and two thermal settings, see Table 1. We refer to the supplementary for details.
Landmarker. We simultaneously train our landmark prediction model using a sparse 70-point 3DA-2D and dense 478-point 2D landmark convention with an image size of $224\times 224$ on our new T-FAKE thermal dataset and the original FAKE RGB dataset. Due to the limited, public availability of the dense 3DA-2D ground truth in the FAKE dataset [61], we use the 478-point 2D predictions of the MediaPipe model for the synthetic RGB images as our dense ground truth dense annotations.
During inference, we use a multi-scale sliding window evaluation to generalize our model to varying image sizes without direct face tracking. For our final landmark prediction, we pool all predictions and use the landmark with the smallest predicted standard deviation across all scales and all sliding windows. We refer to the supplementary for more details.
@@
Evaluation Metrics for Landmarking. We report the normalized mean error (NME) [61, 58, 68] for evaluation, i.e., the mean absolute error (MAE) normalized with an image-dependent constant to account for varying image sizes. Often, normalization is performed by the distance between the outermost eye landmarks (NME IO) [61, 58]. However, the CHARLOTTE benchmark dataset contains many examples with side profile views where only one of the eyes is visible. Therefore, we use the average height and width of the bounding box around the face for normalization [68] (NME W/H). Apart from the NME, we report the failure rates for each of the reference methods. The failure rate is the percentage of frames where no faces could be found. Frames without recognized faces are not included nor weighted in the NMEs. The method of rejection varies for each landmarker. We use vanilla parameters except for MediaPipe [40], where we set the minimal detection confidence to $0.01$. For our method, we evaluate it i) on all frames without failure rate and ii) on all frames where the average uncertainty scores $\sigma_{l}$ of the predicted landmarks $l$, $1\leq l\leq L$, is less than a fixed value $\bar{\sigma}$.
@@
FAKE[61, 62] FAKE 0.1055 0.2675 0.1241 0.2534 0.1887
FAKE* FAKE 0.0933 0.2682 0.1312 0.2348 0.1824
Aachen [33] AACHEN 0.1677 0.2405 0.2942 0.1731 0.2054
TFW [37] TFW 0.0787 0.1249 0.0701 0.1311 0.1012
SF-TL54 U-Net [38] SF-TL54 0.2476 0.3134 0.1757 0.3596 0.2829
SF-TL54 Dlib [38] SF-TL54 0.2159 0.2174 0.1365 0.2741 0.2167
Ours dense ($\bar{\sigma}<\infty$) FAKE + T-FAKE 0.0854 0.1395 0.0764 0.1500 0.1132
Ours ($\bar{\sigma}<\infty$) FAKE + T-FAKE 0.0683 0.1239 0.0645 0.1294 0.0969
Ours ($\bar{\sigma}<6\times 10^{-4}$) FAKE + T-FAKE 0.0671 0.1118 0.0593 0.1179 0.0889
Ours ($\bar{\sigma}<2.2\times 10^{-4}$) FAKE + T-FAKE 0.0661 0.1018 0.0551 0.1059 0.0815
Failure rate (%) $\downarrow$ Aachen [34] AACHEN 49.32 48.72 72.85 25.15 49.01
Mediapipe [40] undisclosed 25.05 46.85 50.09 22.40 36.25
@@
TFW [37] TFW 13.14 21.55 19.26 15.66 17.46
SF-TL54 U-NET / Dlib [38] SF-TL54 29.29 22.57 38.16 13.50 25.84
Ours ($\bar{\sigma}<6\times 10^{-4}$) FAKE + T-FAKE 0.98 10.34 6.94 4.64 5.79
Ours ($\bar{\sigma}<2.2\times 10^{-4}$) FAKE + T-FAKE 6.52 32.76 23.33 16.68 20.00
Table 2: Results on CHARLOTTE dataset splits. Pre-processing with the pre-processing stack for RGB landmarkers is indicated by *. Best NME value is in bold. We denote the confidence threshold with $\bar{\sigma}$. We include all predictions for $\bar{\sigma}<\infty$, Star and 3FabRec resulting in an optimal failure rate of 0 by design. Methods in gray used additional supervision from bounding boxes generated with the ground truth landmarks. RGB datasets are indicated by .
@@ 5.2 Benchmark Datasets for Landmarking
To compare thermal performance, we use the CHARLOTTE dataset due to its high variability in resolution, pose, and background temperature. The dataset is annotated with either 72 2D landmarks for frontal images or 43 2D landmarks for lateral images. The 43-point landmarks represent the visible landmarks in profile pictures with our information on orientation. We use a total of 7684 annotated frames with different resolutions with a minimal one of around 70x60 from the dataset. For evaluation, we propose to split the CHARLOTTE dataset based on image size and perspective. The High condition contains all images larger than the median image width of 200 pixels and Low all smaller images. The condition Side includes all images with profile annotations (43 landmarks), and Front contains all frontal faces (73 landmarks). In addition, we evaluate on the established RGB landmark dataset 300W [58] to verify RGB landmark prediction performance. It uses the classical 68-point convention. Here, we crop images to the pre-defined bounding boxes around each face for our evaluation. We follow the protocol from [65] for 300W evaluation. The average image resolution is about 290 pixels.
300W (NME IO $\downarrow$)
Method Full Comm. Chal.
Mediapipe [40] 0.0562 0.0469 0.0947
FA [6] 0.0435 0.0380 0.0661
Star [65] 0.0287 0.0252 0.0430
Ours 0.0522 0.0468 0.0743
Table 3: Comparison of RGB performance of our method. We trained the label adaptation on the 300W train split and evaluated each method on the 300W test split [24, 65]. For Star, we use the authors’ model pre-trained on 300W. NME has been normalized with respect to interocular distance [61, 58]. Methods with label adaptation are put in italic.
Model FID $\downarrow$ TUFTS MSE [48] $\downarrow$ DRIVE-IN MSE $\downarrow$ [13]
@@
Pix2Pix [26] .2003 .0421 .1453
CycleGAN [66] .2228 .0521 .0730
CUT [49] .2140 .0468 .1370
QS-Attn [23] .1173 .0409 .1303
Table 4: Influence of the regularizers on the thermalization and comparison to other models. The parameter $\lambda_{W}$ corresponds to the impact of the patch-based regularizer (Eq. (2)) and $\lambda_{R}$ corresponds to the segmentation-based regularizer (Eq. (3)). We use $C=(5\cdot 8^{2})^{(-1)}$. The best value is in bold and the second-best is underlined.
@@ 5.3 Comparison for Landmarking
We compare the landmarker performance with training on our full synthetic datasets against state-of-the-art pre-trained models on the CHARLOTTE dataset. Besides relevant thermal landmarkers we include RGB landmarkers. Performance of RGB landmarkers on thermal images can be boosted by applying a stack of unlearned linear filters before prediction [13]. Consequently, application of this pre-processing stack allows us to include RGB landmarkers in our benchmark and to draw a fair comparison beyond task-specific thermal landmarkers. We use label adaptation for all methods to account for different landmark conventions for the image evaluation. We train each label adaptation model with a 1000-image training split on CHARLOTTE similar to [61] and exclude the training split from the final evaluation. Note that this approach allows us to include our sparse model and our dense model in the evaluation. To benchmark the RGB performance of the multimodal model, we evaluate our sparse model predictions against state-of-the-art RGB sparse landmarkers on the 300W dataset.
@@
CHARLOTTE performance. The NME performance as well as the failure rates of the different methods varies drastically for the different CHARLOTTE splits. We list the dataset on which each of the methods has been trained in Table 2. We observe the lowest performance across the landmarkers on the Front condition, followed by the low-resolution images. The FA and Mediapipe RGB methods show a performance that is comparable with the SF-TL54 and Aachen thermal landmarkers. Using a pre-processing stack [13], they consistently outperform them on the High condition in terms of NME and on Full, Front, Low and Side in terms of failure rate. Our model achieves state-of-the-art performance for sparse thermal landmarking across all different conditions. The only model with comparable performance is TFW which, however, only predicts a set of 5 landmarks. While the NME of TFW is slightly lower than our method for $\bar{\sigma}<\infty$ for Low, Front, and Full, our method outperforms TFW on all dataset splits in terms of NME and failure rate for $\bar{\sigma}<6\times 10^{-4}$. These results are presented in Table 2 and examples are visualized in Fig. 5. The 300W-trained RGB methods Star and 3FabRec are close to our method and TFW, however, they have been supplied with the ground truth landmarks for tight bounding box computation. This gives an advantage for label adaptation of broken profile images as well as evaluation on broken and shifted ground truth in CHARLOTTE.
@@
300W performance. Additionally, we evaluate the performance of our multimodal model on the 300W dataset in Table  3. Our model, trained on a combination of thermal and RGB facial images with label adaptation performs between MediaPipe and FA with label adaptation and all methods are outperformed by Star. It is important to note, that the bounding box has only been optimized for Star and our model has been trained without centered faces and a large variability of head positions in the input frames.

### v3

@@ 3.2 The Dataset
Having trained $T_{\theta}$, we use it to thermalize the FAKE dataset which results in the thermal dataset T-FAKE. For each thermal condition, the dataset contains 100k images featuring at least 2k unique subjects. Each image is available in a simulated ‘cold’ and a ‘warm’ environment adding up to 200k images. This number of frames and subjects surpasses the numbers in other thermal landmarking datasets [33, 62, 13, 51, 36, 37, 3] known to the authors. In contrast to thermal landmarking datasets with a fixed distance between the subjects and the camera [33, 51, 37], the distance varies within an estimated range of 80cm to 150cm for the T-FAKE dataset. Moreover, the synthesized images are based on facial scans with a wide range of age and ethnicity and a balanced number of people identifying as male or female [60]. In addition to the original ground truth sparse 3DA-2D landmark annotations, we add dense 2D annotations by transferring Mediapipe [39] predictions for the FAKE to the T-FAKE dataset. This makes our dataset the first densely annotated thermal landmarking dataset. While we found many large thermal landmarking datasets to be non-public [62, 13, 51], our complete dataset with both thermal conditions and annotations is available for download. In combination with the original FAKE dataset, we can train multimodal facial landmarkers jointly on RGB and thermal images as described in the next section. Table 1 displays a comparison between thermal landmarking datasets.
@@ 4 Facial Landmarking
Learning with synthetic images has proven to be very useful in facial landmarking [62, 60, 61]. Based on the proposed T-FAKE and the original FAKE dataset, we can train a supervised thermal or multimodal landmark prediction model for thermal and RGB images using synthetic data only. Moreover, we adapt a model to new annotation conventions by training a small-scale label adaptation model. The data flow for these models is visualized in Figure 3.
@@ 4.1 Landmark Prediction
the authors of [61] proposed to learn the network $T_{\psi}\coloneqq(T_{\psi}^{1},T_{\psi}^{2})$ with $X\mapsto(\boldsymbol{\mu},\boldsymbol{\sigma}^{2})$, where $\boldsymbol{\mu}=(\mu_{l})_{l=1}^{L}\in(\mathbb{R}^{2})^{L}$ and $\boldsymbol{\sigma}^{2}=(\sigma_{l}^{2})_{l=1}^{L}\in\mathbb{R}^{L}$ by minimizing the loss function
$\mathcal{L}(\psi)$ $\coloneqq\sum_{l=1}^{L}\Big{(}\sum_{j=1}^{J}\log T^{2}_{\psi,l}(X^{j})+\frac{\|T^{1}_{\psi,l}(X^{j})-y_{l}^{j}\|_{2}^{2}}{2\,T^{2}_{\psi,l}(X^{j})}\Big{)}.$ (10)
During training, we clip $T^{2}_{\psi,l}\geq\varepsilon>0$ to avoid numerical instability due to too small variances. The first sum controls the location accuracy and the second one the uncertainty prediction. Having learned the network $T_{\psi}$, it produces not only the expected values $\boldsymbol{\mu}$ of the landmarks, but also a measure for its uncertainty $\boldsymbol{\sigma}^{2}$. We threshold the average uncertainty score of multiple sliding windows to enable face detection in our implementation and refer to this as Gaussian log-likelihood with sliding windows (GLL+SW).
@@ 4.2 Label Adaptation
A common evaluation approach that is especially crucial for synthetic data and comparison of 2D and 3DA-2D landmarks is label adaptation [60, 61]. It allows us to adapt to different landmark conventions between the training dataset and an evaluation dataset. Given two sequences $(\hat{\mathbf{y}}_{k},\mathbf{y}_{k})_{k=1}^{K}$ with predicted landmarks $\hat{\mathbf{y}}_{k}\coloneqq(\hat{y}^{l}_{k})_{l=1}^{\hat{L}}\in\mathbb{R}^{2\hat{L}}$, and with ground truth landmarks $\mathbf{y}_{k}\coloneqq(y^{l}_{k})_{l=1}^{L}\in\mathbb{R}^{2\hat{L}}$ with possibly $\hat{L}\neq L$, we aim to train a third label adaptation model $T_{\zeta}:\mathbb{R}^{2\hat{L}}\to\mathbb{R}^{2{L}}$. Here, we use the loss
@@ 5.1 Implementation Details
Thermalization. We train on the RGB+Thermal SEJONG dataset recorded in lab conditions and the large-scale synthetic FAKE dataset featuring in-the-wild RGB images. We train on a resolution of $256\times 256$. The resolution-agnostic U-net allows to apply our two models trained on $256\times 256$ images to the FAKE dataset at the original $512\times 512$ resolution to generate the T-FAKE dataset. This results in the up-to-date largest thermal landmarking dataset with 200k images of at least 2k unique subjects and two thermal settings, see Table 1. We refer to the supplementary for details.
Face-Detection+Landmarker. We simultaneously train our landmark prediction model using a sparse 70-point 3DA-2D and dense 478-point 2D landmark convention with an image size of $224\times 224$ on our new T-FAKE thermal dataset and the original FAKE RGB dataset. Additionally, we mix texture images with random landmarks as negative samples into the training to create meaningful uncertainty estimates $\sigma_{l}$. Due to the limited, public availability of the dense 3DA-2D ground truth in the FAKE dataset [60], we use the 478-point 2D predictions of the MediaPipe model for the synthetic RGB images as our dense ground truth dense annotations.
During inference, we use a multi-scale sliding window evaluation to generalize our model to varying image sizes without direct face tracking beyond pre-defined bounding boxes. For our final landmark prediction, we pool all predictions and use the window with the smallest average standard deviation across all scales and all sliding windows. To integrate face detection into the model, we use the average $\bar{\sigma}$ of the landmark uncertainties $\sigma_{l}$, $1\leq l\leq L$ of the final window per frame and detect faces only if $\bar{\sigma}$ is less than a fixed threshold $t$, i.e., GLL+SW ($\bar{\sigma}<t$). Hence, $\bar{\sigma}<\infty$ corresponds to evaluation on all frames. We refer to the supplementary for more details.
@@
Evaluation Metrics for Landmarking. We report the normalized mean error (NME) [60, 57, 67] for evaluation, i.e., the mean absolute error (MAE) normalized with an image-dependent constant to account for varying image sizes. Often, normalization is performed by the distance between the outermost eye landmarks (NME IO) [60, 57]. However, the CHARLOTTE benchmark dataset contains many examples with side profile views where only one of the eyes is visible. Therefore, we use the average height and width of the bounding box around the face for normalization [67] (NME W/H). Apart from the NME, we report the failure rates for each of the reference methods. The failure rate is the percentage of frames where no faces could be found. Frames without recognized faces are not included nor weighted in the NMEs. The method of rejection varies for each landmarker. We use vanilla parameters except for MediaPipe [39], where we set the minimal detection confidence to $0.01$. We report T-FAKE landmarker results for different levels of face detection confidence to illustrate the trade-off between failure rate and NME.
@@
DAN [32] AACHEN 0.1677 0.2405 0.2942 0.1731 0.2054
YOLO5Face [36] TFW 0.0787 0.1249 0.0701 0.1311 0.1012
U-Net+Wing [37] SF-TL54 0.2476 0.3134 0.1757 0.3596 0.2829
Dlib [37] SF-TL54 0.2159 0.2174 0.1365 0.2741 0.2167
GLL+SW ($\bar{\sigma}<\infty$) [60, 61] FAKE 0.1055 0.2675 0.1241 0.2534 0.1887
GLL+SW ($\bar{\sigma}<\infty$) FAKE 0.0933 0.2682 0.1312 0.2348 0.1824
GLL+SW Dense ($\bar{\sigma}<\infty$) FAKE + T-FAKE 0.0854 0.1395 0.0764 0.1500 0.1132
GLL+SW Sparse ($\bar{\sigma}<\infty$) FAKE + T-FAKE 0.0683 0.1239 0.0645 0.1294 0.0969
GLL+SW Sparse ($\bar{\sigma}<6\times 10^{-4}$) FAKE + T-FAKE 0.0671 0.1118 0.0593 0.1179 0.0889
GLL+SW Sparse ($\bar{\sigma}<2.2\times 10^{-4}$) FAKE + T-FAKE 0.0661 0.1018 0.0551 0.1059 0.0815
Failure rate (%) $\downarrow$ Mediapipe [39] undisclosed 25.05 46.85 50.09 22.40 36.25
@@
DAN [33] AACHEN 49.32 48.72 72.85 25.15 49.01
YOLO5Face [36] TFW 13.14 21.55 19.26 15.66 17.46
U-Net+Wing / Dlib [37] SF-TL54 29.29 22.57 38.16 13.50 25.84
GLL+SW ($\bar{\sigma}<6\times 10^{-4}$) FAKE + T-FAKE 0.98 10.34 6.94 4.64 5.79
GLL+SW ($\bar{\sigma}<2.2\times 10^{-4}$) FAKE + T-FAKE 6.52 32.76 23.33 16.68 20.00
Table 2: Results on CHARLOTTE dataset splits. The pre-processing stack for RGB landmarkers [13] is indicated by *. Best NME value is in bold. Failure rate of GLL+SW depends on the freely chosen confidence threshold $\bar{\sigma}$. We include all predictions for $\bar{\sigma}<\infty$, Star and 3FabRec resulting in an optimal failure rate of 0 by design. Methods in gray perform no face detection and are thus reported with additional supervision from bounding boxes generated with the ground truth landmarks. RGB datasets are indicated by .
@@ 5.2 Benchmark Datasets for Landmarking
To compare thermal performance, we use the CHARLOTTE dataset due to its high variability in resolution, pose, and background temperature (see Figure 4). The dataset is annotated with either 72 2D landmarks for frontal images or 43 2D landmarks for lateral images. The 43-point landmarks represent the visible landmarks in profile pictures with our information on orientation. We use a total of 7684 annotated frames with different resolutions with a minimal one of around 70x60 from the dataset. For evaluation, we propose to split the CHARLOTTE dataset based on image size and perspective. The High condition contains all images larger than the median image width of 200 pixels and Low all smaller images. The condition Side includes all images with profile annotations (43 landmarks), and Front contains all frontal faces (73 landmarks). In addition, we evaluate on the established RGB landmark dataset 300W [57] to verify RGB landmark prediction performance. It uses the classical 68-point convention. Here, we crop images to the pre-defined bounding boxes around each face for our evaluation. We follow the protocol from [64] for 300W evaluation. Average image resolution is about 290 pixels.
Model FID $\downarrow$ TUFTS MSE [47] $\downarrow$ DRIVE-IN MSE [13] $\downarrow$
Pix2Pix [25] .2003 .0421 .1453
CycleGAN [65] .2228 .0521 .0730
CUT [48] .2140 .0468 .1370
QS-Attn [22] .1173 .0409 .1303
@@
Table 3: Comparison to other models and regularizer influence on the thermalization. The parameter $\lambda_{W}$ corresponds to the impact of the patch-based regularizer (Eq. (2)) and $\lambda_{R}$ corresponds to the segmentation-based regularizer (Eq. (3)). We use $C=(5\cdot 8^{2})^{-1}$. The best value is in bold and the second-best is underlined.
@@ 5.3 Comparison for Landmarking
We compare the landmarker performance with training on our full synthetic datasets against pre-trained models on the CHARLOTTE dataset. Besides relevant thermal landmarkers we include state-of-the-art RGB landmarkers. Thermal performance of RGB landmarkers can be boosted by applying a stack of unlearned linear filters before prediction [13]. Consequently, such pre-processing allows for a fairer comparison between task-specific thermal landmarkers and RGB landmarkers. We use label adaptation for all methods to account for different landmark conventions for the image evaluation. We train each label adaptation model with a 1000-image training split on CHARLOTTE similar to [60] and exclude the training split from the final evaluation. Note that this approach allows us to include our sparse model and our dense model in the evaluation. To benchmark the RGB performance of the multimodal model, we evaluate our sparse model predictions against state-of-the-art RGB sparse landmarkers on the 300W dataset.
@@
CHARLOTTE performance. The NMEs and the failure rates of the different landmarkers vary drastically for the CHARLOTTE splits with generally worse performance for the Front and Low conditions. As for the landmarkers, we observe a clear advantage of GLL+SW ($\bar{\sigma}<\infty$) trained on FAKE+T-FAKE over GLL+SW ($\bar{\sigma}<\infty$) trained on FAKE only across sparse and dense landmarks, metrics, and splits. This confirms the suitability of our T-FAKE dataset for thermal landmarking. The FA and Mediapipe RGB landmarkers and the SF-TL54 and Aachen thermal landmarkers show comparable performance. Using a pre-processing stack [13], leads to a drastic failure reduction for the RGB landmarkers. Our T-FAKE model achieves state-of-the-art performance for sparse thermal landmarking across all different conditions. The only models with comparable performance are Star which, however, performs no face detection and is evaluated with ground truth bounding boxes, and the TFW model which, however, only predicts a set of 5 landmarks. While the NME of the TFW model is slightly lower than our model for $\bar{\sigma}<\infty$ for Low, Front, and Full, our model outperforms Star and TFW on almost all dataset splits in terms of NME and failure rate for $\bar{\sigma}<6\times 10^{-4}$. These results are presented in Table 2 and examples are visualized in Figure 4. The 300W-trained RGB methods Star and 3FabRec use ground truth information for tight bounding box computation. Integration of face detection into their evaluation would presumably decrease the reported NMEs.
@@
300W performance. We evaluate the RGB performance of our multimodal model on the 300W dataset in Table  4. Our model, trained on a combination of thermal and RGB facial images, with label adaptation performs between MediaPipe and FA with label adaptation. All methods are outperformed by Star. Note that Star is the only method trained on 300W landmarks and bounding boxes. Our implementation is the only one trained on general facial images and not on bounding boxes. Most importantly, training with FAKE only does not improve RGB performance.
300W (NME IO $\downarrow$)
Method Training Data Full Comm. Chal.
Star [64] 300W 0.0287 0.0252 0.0430
Mediapipe [39] undisclosed 0.0562 0.0469 0.0947
FA [6] LS3D-W 0.0435 0.0380 0.0661
GLL+SW ($\bar{\sigma}<\infty$) FAKE 0.0570 0.0494 0.0874
GLL+SW ($\bar{\sigma}<\infty$) FAKE + T-FAKE 0.0522 0.0468 0.0743
Table 4: Comparison of RGB performance. Label adaptation is trained on the 300W train split. Each method is evaluated on the 300W test split [23, 64]. For Star, we use the authors’ model pre-trained on 300W. NME has been normalized with the interocular distance [60, 57]. Methods with label adaptation are put in italic. Note that training with FAKE + T-FAKE does not degrade performance on 300W compared to FAKE alone.